OpenStudy (anonymous):

Find the point on the curve y=ln(x^2) where the slope of the tangent line is 2/3 find the x intercept of the line tangent to the curve at that point

zepdrix (zepdrix):

\[\Large\bf\sf y(x)\quad=\quad \ln(x^2)\]And we want to find the x value that gives us,\[\Large\bf\sf y'(x)\quad=\quad \frac{2}{3}\]

zepdrix (zepdrix):

Do you know how to take the derivative of the function?

OpenStudy (anonymous):

yes it is 2/x i believe

OpenStudy (anonymous):

omy do i just plug in :o

zepdrix (zepdrix):

Mmm ok cool. \[\Large\bf\sf y'(x)\quad=\quad \frac{2}{3}\]Setting our derivative equal to 2/3,\[\Large\bf\sf \frac{2}{x}\quad=\quad \frac{2}{3}\]

zepdrix (zepdrix):

Yah the first part isn't too bad :) Find the x value where this is true.

OpenStudy (anonymous):

so x =3

zepdrix (zepdrix):

Mmmmk good.

OpenStudy (anonymous):

so that is my point i assume ?

zepdrix (zepdrix):

Hmm I'm not sure what the fastest method for finding the x-intercept is... This is the process I took though, We find an equation for the tangent line at x=3. \[\Large\bf\sf y_{\tan}\quad=\quad \frac{2}{3}x+b\] After we find the y-intercept we can plug in y=0 to find the x-intercept. If there is a faster way you can stop me Kanui :P hehe

OpenStudy (anonymous):

well i've learned finding the x you just set the y to zero or if you want to find y you make x 0 but I forgot if thats intercept. :(

zepdrix (zepdrix):

Yah that's the right process :) But what y are we setting equal to zero? We don't have an equation for the tangent line yet. After we set it up we can set y=0 to find the x-intercept of that line.

zepdrix (zepdrix):

We could do this in point-slope form if you prefer, but slope-intercept seems fine :p We need a coordinate pair for the line. So we'll plug x=3 into our function to find a corresponding y value.

zepdrix (zepdrix):

\[\Large\bf\sf y(3)\quad=\quad \ln(3^2)\quad=\quad 2\ln(3)\]So our coordinate pair is,\[\Large\bf\sf (3,\;2 \ln3)\]

zepdrix (zepdrix):

Umm actually let's do point-slope form. It'll make more sense I think.\[\Large\bf\sf y-y_o\quad=\quad m(x-x_o)\]Plugging in our slope, and our coordinate point,\[\Large\bf\sf y-2\ln3\quad=\quad \frac{2}{3}(x-3)\]This is the equation of our line tangent to the curve at x=3. (which has slope 2/3 as they requested in the first part). Now to find our x-intercept we let y=0.

OpenStudy (anonymous):

im a little confused

zepdrix (zepdrix):

Ya it's a lot going on :(

OpenStudy (anonymous):

wait do i only look at point slope form part?

zepdrix (zepdrix):

We're looking for the x-intercept of our `tangent line`. Before we could find the x-intercept we had to establish the equation for the tangent line. Whether you write the tangent line in `point-slope` form or `slope-intercept` form doesn't matter. Which form are you more familiar with? They're both just different ways of writing linear equations.

OpenStudy (anonymous):

i think i understand the way it is written in point slope the best

zepdrix (zepdrix):

So then hopefully this looks a little bit familiar:\[\Large\bf\sf y-y_o\quad=\quad m(x-x_o)\] We have to plug in a coordinate pair \(\Large\bf\sf (x_o,y_o)\) and a slope \(\Large\bf\sf m\).

zepdrix (zepdrix):

The slope was `given to us` at the start of the problem. We just had to figure out `where that slope was located`.

OpenStudy (anonymous):

so it could be y-yo= 2/3(x-xo) or am i totally off

OpenStudy (anonymous):

the Y im not exactly sure where to get

zepdrix (zepdrix):

Plugging in the slope they gave us? yes good.

zepdrix (zepdrix):

We finished the first part, determining that the `location` of the 2/3 slope is at x=3 right?

OpenStudy (anonymous):

sorrry im like eyes half open so i didnt even actually plug in 3 :O

OpenStudy (anonymous):

but yes we just determined that

zepdrix (zepdrix):

So we need the y coordinate that corresponds to x=3. So we plug the x=3 into the `original function`. brb gotta get my food out the oven :OOO

OpenStudy (anonymous):

i got 2.19. thank you for your help I think thats the farthest I can go I really need to sleep and shower ! thank you again I appreciate it so much my grade depends on this assignment. I wish i loved math ahha goodnight mr/mrs zepdrix

zepdrix (zepdrix):

lol night :D

zepdrix (zepdrix):

(3, 2.19) Plug it into the x0, y0 in the point-slope form. then y=0 gives you your x value. go to bed. :P